[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Jul 24, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Jul 25 02:56:29 UTC 2017


HSU Shift Log for Mon Jul 24 2017 22:56:23 GMT-0400 (EDT)

Shift: R

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: Coordinated with Acting HSU BC, who arranged with SDM  
to produce GOES-16 Meso Window over H. Hilary on Tuesday, 25 July 2017,  
from 0900Z to 1500Z.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170725/f046b8fc/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list